# Diff-Control: A Stateful Diffusion-based Policy for Imitation Learning

### v2

@@ Abstract
While imitation learning provides a simple and effective framework for policy learning, acquiring consistent action during robot execution remains a challenging task. Existing approaches primarily focus on either modifying the action representation at data curation stage or altering the model itself, both of which do not fully address the scalability of consistent action generation. To overcome this limitation, we introduce the Diff-Control policy, which utilizes a diffusion-based model to learn action representation from a state-space modeling viewpoint. We demonstrate that diffusion-based policies can acquire statefulness through a Bayesian formulation facilitated by ControlNet, leading to improved robustness and success rates. Our experimental results demonstrate the significance of incorporating action statefulness in policy learning, where Diff-Control shows improved performance across various tasks. Specifically, Diff-Control achieves an average success rate of 72% and 84% on stateful and dynamic tasks, respectively. Notably, Diff-Control also shows consistent performance in the presence of perturbations, outperforming other state-of-the-art methods that falter under similar conditions. Project page: https://diff-control.github.io/
@@ I Introduction
In the field of robotics, imitation learning [1] has become a prominent method for programming robots based on expert demonstrations in a data-driven and sample-efficient manner. Within this context, behavioral cloning, in its simplest form, is structured as a supervised regression task aimed at mapping observations to corresponding actions. Previous studies have explored various approaches to learning behavioral cloning policies, such as directly outputting actions via regression models [2] or utilizing implicit policies [3]. Notably, diffusion-based policies [4] have emerged as a standout choice due to their ability to model multimodal action distributions effectively, leading to enhanced performance.
In practice, however, the concern over inconsistency in action representation remains a persistent challenge. Such inconsistencies can lead to noticeable disparities between the distribution of robot trajectories and the underlying environment, thereby limiting the efficacy of control policies [5]. The primary causes of this inconsistency typically stem from the context-rich nature of human demonstrations [6], distribution shift problems [7], and the volatile nature of high-dynamic environments. Previous approaches, such as action chunking [8] and predicting closed-loop action sequences [4], have been proposed to address this issue. Additionally, Hydra [9] and Waypoint-based manipulation [10] modify action representations to ensure consistency. However, these approaches address the problem by altering the action representation without using the actions as is.
Instead, can we learn to explicitly impose temporal consistency by incorporating temporal transitions within diffusion policies? In the realm of deep state-space models [11, 12, 13], the effective learning of a state transition model enables the identification of underlying dynamic patterns. In this paper, we argue that such deep transition models can easily be integrated into diffusion policies in order to capture temporal action dynamics as a state-space model. This integration makes the policy stateful, thereby increasing robustness and success rates.
@@
Figure 2: The Diff-Control Policy, denoted as $\pi_{\boldsymbol{\theta}}$, is a stateful diffusion policy. It operates by generating a sequence of actions while incorporating conditioning on previously generated actions. In this example, the Diff-Control policy is depicted executing the “Open Lid” task. For instance, in the second sub-figure, the blue trajectory represents previous action trajectory, denoted as ${\bf a}_{[W_{t}]}$, while the red trajectory displays the newly generated sequence of actions, denoted as ${\bf a}_{[W_{t+h}]}$.
@@
  • A Bayesian formulization within diffusion-based polices using ControlNet structure as a transition model to ensure consistent action generation.
  • Diff-Control stateful policy representation performing dynamic and temporal sensitive tasks with at least 10% and 48% improvement in success rate. Diff-Control policy exhibits notable precision and robustness against perturbations, achieving at least a 30% higher success rate compared to state-of-the-art methods.
  • Empirical evidence that Diff-Control Policy can perform on a wide range of tasks including high-precision tasks with at least 31% success rate improvement.
II Related work
Diffusion models have gained considerable attention in diverse robot control applications, including tasks such as planning [15], motion prediction [16] and reinforcement learning [17]. Notably, diffusion-based control policies have garnered substantial interest as a robust representation of agent behavior [4, 15]. In the field of imitation learning, these diffusion-based policies have been applied across various tasks, leveraging their capacity to articulate multimodal action distributions. Illustrative applications include goal-directed navigation for mobile robots [18], integration of human behavior in collaborative human-robot tasks [19], and language-guided robot skill learning activities [20]. While these policies demonstrate remarkable scalability in diverse control domains, they are fundamentally stateless, lacking provisions for incorporating memory and prior knowledge into the controller, potentially leading to inconsistent action generation.
To overcome this constraint, integrating a stateful policy [21, 22] becomes imperative, particularly in scenarios where robots are operating in dynamic environments or engaging in tasks over extended horizons. Stateful policies leverage both historical and present states to formulate actions, facilitating adept management of intricate tasks. The application of stateful trajectories has shown benefits in long-horizon skill planning [23], and showcases predictive prowess in evolving environments [24]. A central challenge in this regard is the state representation itself - which features have to be part of the temporal state in order to ensure accurate completion of the task.
In our study, we adopt a deep state-space model (DSSM) approach [11] to address these challenges. DSSMs enable the effective learning of end-to-end models that capture state transitions from observed sequences alone [12, 13]. Differentiable Filters [13] as a subclass of DSSMs has shown the effectiveness of modeling uncertainty with noise profiles in state space modeling. These Differentiable Filters also show good performance on real-world tasks with considerable improvement in state tracking accuracy [25, 26, 27]. In our work, we leverage a formulation at the intersection of recursive Bayesian filtering and deep learning [26, 28] to construct a state-space model that captures the underlying transition dynamics in both space and time.
III Method
@@
Figure 2 offers an overview of Diff-Control in action for the “Open Lid” task. Within each time window (depicted in red), Diff-Control generates action sequences. When generating subsequent action sequences, it utilizes previous actions as an additional control input, shown in blue. This temporal transition is achieved through Bayesian formulation, effectively bridging the gap between standalone policies and state space modeling.
Subsequently, we will introduce the elements of the Diff-Control algorithm, the underlying network architecture and the training process in detail.
III-A Diffusion Model
Diffusion models are generative models that iteratively map Gaussian noise to a target distribution, with the capacity to optionally condition on contextual information [29]. Given ${\bf a}_{T}\sim\mathcal{N}({\bf 0},{\bf I})$ as the starting point, the diffusion model predicts an output sequence as ${\bf a}_{T-1},{\bf a}_{T-2},\cdots,{\bf a}_{0}$, where each subsequent output serves as a denoised version of the previous output. ${\bf a}_{0}$ is the output after the diffusion process. We use the denoising diffusion model (DDPM) [30] as the backbone. In training, the noise inputs can be generated given a varied noise level: ${\bf a}_{\tau}=\sqrt{\bar{\alpha}_{\tau}}{\bf a}+\sqrt{1-\bar{\alpha}_{\tau}}{\bf z}$, where $\bar{\alpha}_{\tau}$ is variance schedule, and ${\bf z}$ is random noise, ${\bf z}\sim\mathcal{N}({\bf 0},{\bf I})$. We can train a neural network $\epsilon(\cdot)$ to predict the noise added to the input by minimizing:
where (${\bf o},{\bf a}$) represents observation and action pairs, $\tau$ is the denoising timesteps, $\tau\in[1,T]$. In the sampling step, we interatively run the denoising process:
${\bf a}_{\tau-1}=\frac{1}{\sqrt{\alpha_{\tau}}}\left({\bf a}_{\tau}-\frac{1-\alpha_{\tau}}{\sqrt{1-\bar{\alpha}_{\tau}}}\epsilon({\bf o},{\bf a}_{\tau},\tau)\right)+\sigma_{\tau}{\bf z},$ (2)
where $\sigma_{\tau}$ and $\alpha_{\tau}$ are noise schedule parameters, which has been well-studies in [30, 31].
III-B Recursive Bayesian Formulation
@@
where $\eta$ is a normalization factor, $p({\bf o}_{t}|{\bf a}_{t},{\bf c})$ is the observation model and $p({\bf a}_{t}|{\bf a}_{t-1},{\bf c})$ is the transition model. The transition model describes the laws that govern the evolution of the system dynamics, while the observation model identifies the relationship between the internal state of the system and observed, noisy measurements.
III-C Diff-Control Policy
We now show how Bayesian formulation and diffusion model can be coupled together such that one policy can generate stateful action sequences that facilitate consistent robot behaviors. We propose Diff-Control Policy $\pi_{\boldsymbol{\theta}}({\bf a}_{[W_{t}]}|{\bf o},{\bf a}_{[W_{t-h}]},{\bf c})$ as shown in Figure 2, which is parameterized by $\boldsymbol{\theta}$. Here, $h$ stands for the execution horizon, ${\bf c}$ represents a language condition in the form of a natural human instruction, and ${\bf o}$ denotes a sequence of images captured by an RGB camera of the scene. The policy $\pi_{\boldsymbol{\theta}}$ generates a window of trajectory ${\bf a}_{[W_{t}]}=[{\bf a}_{1},{\bf a}_{2},\cdots{\bf a}_{W}]^{T}\in\mathbb{R}^{7\times W}$, where $W$ refers to the window size or the prediction horizon.
@@ III-C Diff-Control Policy
Base Policy: To begin, we train a diffusion-based policy [4] following the step in Section III-A as the base policy $\bar{\pi}_{\boldsymbol{\psi}}({\bf a}_{[W_{t}]}|{\bf o},{\bf c})$. We adopt the 1D temporal convolutional networks from [15] and construct the U-net backbone. The policy $\bar{\pi}_{\boldsymbol{\psi}}$ can autonomously execute and generate actions without any temporal information dependency.
@@
III-D Training
The training process for the base policy $\bar{\pi}_{\boldsymbol{\psi}}({\bf a}_{[W_{t}]}|{\bf o},{\bf c})$ follows a straightforward approach. We firstly encode the observation ${\bf o}$ and language condition ${\bf c}$ into the same embedding dimension, and then we utilize the learning objective defined in Equation 1 to train the base policy. The same learning objective is used in finetuning the ControlNet:
@@ III-D Training
III-E Design Decisions
For all our experiments, we adopt a CNN-based U-net architecture. This choice of a CNN-based backbone has proven to be suitable for a diverse range of tasks, both in simulated and real-world scenarios. In our model architecture, as depicted in Figure 4, the encoder and decoder blocks utilize 1D convolutional layers with varying kernel sizes. To implement the zero convolutional layers for ControlNet, we employ 1D $1\times 1$ convolutional layers with weights initialized to zero. This approach ensures that any potential detrimental noise does not affect the hidden states of the trainable neural network layers during the initial stages of training [14].
We use a window size $W=24$ as the default predicting horizon for all the experiments, and the execution horizon $h$ between ${\bf a}_{[W_{t-h}]}$ and ${\bf a}_{[W_{t}]}$ are 12 steps. We only execute 12 steps ($h=12$) by default during the experiments in align with [4]. Because according to [4], the execution horizon being too large or too small can cause the drop of the performance.
During the implementation of language-conditioned tasks, we observe that diffusion-based policy reaches a certain capacity when learning varied actions utilizing complex CLIP language features [34] as conditions. To address this, a more practical approach is to incorporate a fuse layer and increasing embedding size for visual and language representations, instead of concatenating them directly. This modification can enhance the policy’s overall performance for language-conditioned tasks.
IV Real-Robot Tasks
We conducted a comprehensive evaluation of the Diff-Control Policy by comparing it with four baseline methods across five distinct robot tasks. Table I provides a summary of the task properties. The tasks encompassed in our evaluation include: (a) Language Conditioned kitchen tasks, (b) Open Lid task in the kitchen scene as a high-precision task, (c) Duck Scooping task in a dynamic scenario, (d) Drum Beats as a periodic task.
Figure 5: Real-world tasks in this study: a) language-conditioned pick and place task in kitchen scenario, b) “Open Lid” task with high precision requirement, c) “Duck Scooping” task in a water tank, d) “Drum Beats” task by hitting the drum 3 times.
TABLE I: Task Properties
@@ IV Real-Robot Tasks
Figure 6: Diff-Control for real-world tasks: The top two rows depict the language-conditioned kitchen task, where the Diff-Control Policy successfully generates actions in alignment with the given language command and consistently executes these actions. The third row shows a successful duck scooping experiment. The last row displays one drum task result. The results are best appreciated with videos on the website: https://diff-control.github.io/.
@@
The action of the UR5 robot arm is represented as ${\bf a}_{[W_{t}]}$, where each action is denoted as ${\bf a}_{i}=[x,y,z,\alpha,\beta,\gamma]^{T}$, where $i\in[1,W]$. It encompasses the position of the end-effector in Cartesian coordinates $(x,y,z)$, the orientations $(\alpha,\beta,\gamma)$, and the gripper’s joint angle $g$. For all the tasks, the input modalities consist of two modalities: ${\bf o}$ and ${\bf c}$. The first modality, ${\bf o}\in\mathbb{R}^{224\times 224\times 3}$, corresponds to a RGB image. The second modality, $\bf c$, refers to a language embedding derived from natural language sequences. This embedding serves as the linguistic input for the robot’s understanding and decision-making processes. Table I further arranges the tasks in ascending order of subjective difficulty, providing a summary of task characteristics such as the number of distractors (Dis), number of expert demonstrations (Dem), number of varied actions (Act), and whether high-precision (HiPrec) is required or not. In the following sections, we offer a comprehensive overview of the experimental setups for each task, outline the data collection procedures employed, and the specific challenges encountered within each scenario.
Language Conditioned Kitchen task: This task is designed to resemble several tasks in the kitchen scenario [35, 36]. The robot workspace consists of a scaled-down real-world model kitchen, as illustrated in Figure 5(a). The kitchen environment encompasses various objects, including pots, pans, bowls, and distractor objects resembling plush vegetables. During the data collection process, the distractor objects are randomly positioned. Trained experts are responsible for teleoperating the robot to perform two specific actions within the kitchen environment. The actions involve fetching a tomato and subsequently placing it either in the pot on the stove (A) or in the sink (B) based on the given language instruction.
High-Precision Open Lid: The task includes lifting the lid and subsequently placing it onto a nearby bowl, necessitating precise control. Depicted in Figure 5(b), the lid handle is relatively small, and the lid surface is reflective. To gather data for this task, we obtain 50 expert demonstrations. For each demonstration, we introduce random placements of 5 or more distractor objects, as well as slight shifts in the pot’s position and rotations of the lid.
@@
Duck Scooping: Inspired by [37], we explore the interaction between the policy and fluid dynamics. In this task, we equip the robot with a ladle and the robot’s objective is to scoop the duck out of the water. As depicted in the bottom right of Figure 5(c), this task presents challenges due to perturbations caused by the entry of the ladle into the water. The flow of water affects the position of the rubber duck, necessitating the robot to execute precise and cautious movements in order to successfully capture the duck.
@@
Drum Beats: This task is specifically designed for robots to learn periodic motions, a challenging feat due to the unique action representation required [38]. As illustrated in Figure 5(d), the task presents difficulty as the robot must accurately count the number of drum beats and determine when to cease drumming. A total of 150 expert demonstrations were obtained by teleoperating the robot to strike the drum three times in each demonstration.
@@
Method Kitchen Lang. Kitchen Lid Duck Scoop Drum Beats
Pick & Place Duration Open Lid Duration Scoop Duration Hit $\times$ 3 Duration
Image-BC 60% 34.48$\pm$1.46 12% 35.13$\pm$0.50 16% 21.04$\pm$1.99 0% 19.50$\pm$0.79
ModAttn 28% 36.04$\pm$1.51 36% 46.42$\pm$4.69 12% 20.43$\pm$0.16 0% 25.74$\pm$0.68
BC-Z LSTM 64% 35.78$\pm$1.69 40% 46.33$\pm$7.20 36% 21.21$\pm$1.54 0% 17.53$\pm$0.31
Diffusion Policy 88% 41.23$\pm$2.12 64% 47.94$\pm$1.95 76% 21.72$\pm$2.17 24% 20.60$\pm$1.64
@@
V Evaluation
Figure 7: Left shows Diff-Control successfully recovers from a human perturbation, efficiently picking up the tomato; Right shows when visual occlusion is applied, Diff-Control manages to scoop the duck from the water successfully. In contrast, the diffusion policy fails. Red shade implies perturbation.
The efficacy of the proposed policy is evaluated through four experiments as described in Section IV. These experiments aim to address the following questions: (a) Can the Diff-Control policy demonstrate generalization capabilities across diverse tasks? (b) To what extent does the Diff-Control policy outperform the current state-of-the-art methods in terms of overall performance? (c) What are the distinguishing characteristics and benefits of utilizing a stateful policy compared to a non-stateful policy?
We propose the baselines are as follows:
  1. 1.
    Image-BC: This baseline adopts an image-to-action agent framework, similar to BC-Z [2], it is built upon ResNet-18 backbone and employs FiLM [39] for conditioning using CLIP language features.
  2. 2.
    ModAttn [32]: This method employs a transformer-style neural network and uses a modular structure to address each sub-aspects of the task via neural attention, it requires a human expert correctly identifies components and subtasks to each task.
  3. 3.
    BC-Z LSTM: This baseline represents a stateful policy inspired by the BC-Z architecture. The incorporation of a prior input is achieved by fusing the prior actions and language conditions using MLP and LSTM layers.
  4. 4.
    Diffusion Policy [4]: This baseline is a standard diffusion policy.
For all experiments, we present the results obtained from the best performing configuration of each baseline method. All the baseline models are reproduced and trained using the same collected expert demonstrations for a total of 3,000 epochs. Throughout the training process, checkpoints are saved every 300 epochs. In our analysis, we report the best results achieved from these saved checkpoints for each baseline method. Each experiment was carried out on batch size of 64 on a single NVIDIA Quadro RTX 8000 GPU for roughly 24 hours. For all the tasks, we use the Adamw optimizer with a learning rate of 1e-4.
V-A Kitchen Task Evaluation
Language Conditioned: The evaluation results for the language-conditioned kitchen task are presented in Table II. The duration of the task is measured from the initiation of running the policy until the robot successfully opens the gripper and places the tomato either in the sink or on the kitchen stove. We tested each policy for 25 trials. In each of these 25 test trials, we randomized the distractor locations and the pot and bowl placements. Figure 6 shows the Diff-Control policy performing the task with two language-conditions. Among all the baseline methods, Diff-Control achieves the highest success rate with 92%, which is 5%, 64%, and 32% higher than diffusion policy, ModAttn, and Image-BC, respectively. All policies are capable of reaching the correct destination based on the language inputs. Interestingly, we observe that Diff-Control policy exhibits the ability to recover from perturbations as shown in Figure 7(left). Diff-Control continues following the language commands whereas ModAttn, and Image-BC failed to do so.
Open Lid: Similar to language-conditioned kitchen task, we randomized the distractor locations and the pot and bowl placements through out the test trials. Diff-Control policy is able to achieve this high-precision task with 80% success rate over 25 trials. The efficacy of the policy is visually illustrated in Figure 2, presenting the diffusion steps involved. In comparison to the baselines, the diffusion policy resulted in a success rate of 64% and the ModAttn policy with 36%. The Image-BC policy yielded the lowest success rate of 0%. It was observed that even a minor offset in the lid position presented substantial challenges for these explicit policies, i.e, Image-BC, ModAttn, and BCZ-LSTM. In contrast, Diff-Control policy demonstrated superior performance and does not show the tendency to overfit on idle actions. This characteristic enabled the gripper to maneuver and explore varied locations for successfully gripping the lid handle.
V-B Duck Scooping Evaluation

### v3

# Enabling Stateful Behaviors for Diffusion-based Policy Learning

@@ Abstract
While imitation learning provides a simple and effective framework for policy learning, acquiring consistent actions during robot execution remains a challenging task. Existing approaches primarily focus on either modifying the action representation at data curation stage or altering the model itself, both of which do not fully address the scalability of consistent action generation. To overcome this limitation, we introduce the Diff-Control policy, which utilizes a diffusion-based model to learn the action representation from a state-space modeling viewpoint. We demonstrate that we can reduce diffusion-based policies’ uncertainty by making it stateful through a Bayesian formulation facilitated by ControlNet, leading to improved robustness and success rates. Our experimental results demonstrate the significance of incorporating action statefulness in policy learning, where Diff-Control shows improved performance across various tasks. Specifically, Diff-Control achieves an average success rate of 72% and 84% on stateful and dynamic tasks, respectively. Project page: https://github.com/ir-lab/Diff-Control
@@ I Introduction
Previous studies have explored various approaches to learning behavioral cloning policies, such as directly outputting actions via regression models [1] or utilizing implicit policies [2]. Notably, diffusion-based policies [3] have emerged as a standout choice due to their ability to model multimodal action distributions effectively, leading to enhanced performance.
In practice however, the concern over inconsistency in action representation remains a persistent challenge. Such inconsistencies can lead to noticeable disparities between the distribution of robot trajectories and the underlying environment, thereby limiting the efficacy of control policies [4]. The primary causes of this inconsistency typically stem from the context-rich nature of human demonstrations [5], distribution shift problems [6], and the volatile nature of high-dynamic environments. Previous approaches, such as action chunking [7] and predicting closed-loop action sequences [3], have been proposed to address this issue. Additionally, Hydra [8] and Waypoint-based manipulation [9] modify action representations to ensure consistency. However, these approaches address the problem by altering the action representation without using the actions as is.
Instead, can we learn to explicitly impose temporal consistency by incorporating temporal transitions within diffusion policies? In the realm of deep state-space models (DSSMs) [10, 11, 12], the effective learning of a state transition model enables the identification of underlying dynamic patterns. In this paper, we argue that such deep transition models can easily be integrated into diffusion policies in order to capture temporal action dynamics as a state-space model. This integration makes the policy stateful, thereby increasing robustness and success rates.
@@
  • A deep, recursive Bayesian filter within diffusion-based polices using ControlNet structure as a transition model to ensure consistent action generation.
  • Diff-Control stateful policy representation performing dynamic and temporal sensitive tasks with at least 8% and 48% improvement in success rate.
  • Diff-Control policy exhibits notable precision and robustness against perturbations, achieving at least a 30% higher success rate compared to state-of-the-art methods.
II Method
@@
II-A Recursive Bayesian Formulation
@@
where $\eta$ is a normalization factor, $p({\bf o}_{t}|{\bf a}_{t},{\bf c})$ is the observation model and $p({\bf a}_{t}|{\bf a}_{t-1},{\bf c})$ is the transition model.
II-B Diff-Control Policy
We now show how Bayesian formulation and diffusion model can be coupled together such that one policy can generate stateful action sequences that facilitate consistent robot behaviors. We propose Diff-Control Policy $\pi_{\boldsymbol{\theta}}({\bf a}_{[W_{t}]}|{\bf o},{\bf a}_{[W_{t-h}]},{\bf c})$, which is parameterized by $\boldsymbol{\theta}$. Here, $h$ stands for the execution horizon, ${\bf c}$ represents a language condition in the form of a natural human instruction, and ${\bf o}$ denotes a sequence of images captured by an RGB camera of the scene. The policy $\pi_{\boldsymbol{\theta}}$ generates a window of trajectory ${\bf a}_{[W_{t}]}=[{\bf a}_{1},{\bf a}_{2},\cdots{\bf a}_{W}]^{T}\in\mathbb{R}^{7\times W}$, where $W$ refers to the window size or the prediction horizon.
@@ II-B Diff-Control Policy
Base Policy: To begin, we train a diffusion-based policy [3] as the base policy $\bar{\pi}_{\boldsymbol{\psi}}({\bf a}_{[W_{t}]}|{\bf o},{\bf c})$. We adopt the 1D temporal convolutional networks from [15] and construct the U-net backbone. The policy $\bar{\pi}_{\boldsymbol{\psi}}$ can autonomously execute and generate actions without any temporal information dependency.
@@
II-C Training
The training process for the base policy $\bar{\pi}_{\boldsymbol{\psi}}({\bf a}_{[W_{t}]}|{\bf o},{\bf c})$ follows a straightforward approach. We firstly encode the observation ${\bf o}$ and language condition ${\bf c}$ into the same embedding dimension, and then we utilize the learning objective defined in Equation 3 to train the base policy.
The same learning objective is used in finetuning the ControlNet:
@@ II-C Training
III Real-Robot Tasks
We conducted a comprehensive evaluation of the ControlNet Policy by comparing it with four baseline methods across two robot tasks. Table I provides a summary of the task properties. The tasks encompassed in our evaluation include: (a) Duck Scooping task in a dynamic scenario, (b) Drum Beats as a periodic task.
Figure 4: Real-world tasks in this study: a) “Duck Scooping” task in a water tank, b) “Drum Beats” task by hitting the drum 3 times.
TABLE I: Task Properties.
@@ III Real-Robot Tasks
Figure 5: Diff-Control for real-world tasks: The first row shows a successful duck scooping experiment. The second row displays one drum task result. The results are best appreciated with videos on the website: https://diff-control.github.io/.
@@
The action of the UR5 robot arm is represented as ${\bf a}_{[W_{t}]}$, where each action is denoted as ${\bf a}_{i}=[x,y,z,r,p,y,g]^{T}$, where $i\in[1,W]$. It encompasses the position of the end-effector in Cartesian coordinates $(x,y,z)$, the orientations $(r,p,y)$, and the gripper’s joint angle $g$. For all the tasks, the input modalities consist of two modalities: ${\bf o}\in\mathbb{R}^{224\times 224\times 3}$, corresponds to a RGB image. ${\bf c}$ refers to a language embedding derived from natural language sequences. Task Properties in Table I includes distractors (Dis), number of demonstrations (Dem), number of actions (Act), and whether high-precision is required (HiPrec).
@@
Duck Scooping: Inspired by [17], we explore the interaction between the policy and fluid dynamics. In this task, we equip the robot with a ladle and the robot’s objective is to scoop the duck out of the water. As depicted in the bottom right of Figure 4(a), this task presents challenges due to perturbations caused by the entry of the ladle into the water. The flow of water affects the position of the rubber duck, necessitating the robot to execute precise and cautious movements in order to successfully capture the duck.
@@
Drum Beats: This task is specifically designed for robots to learn periodic motions, a challenging feat due to the unique action representation required [18]. As illustrated in Figure 4(b), the task presents difficulty as the robot must accurately count the number of drum beats and determine when to cease drumming. A total of 150 expert demonstrations were obtained by teleoperating the robot to strike the drum three times in each demonstration.
IV Evaluation
The efficacy of the proposed policy is evaluated through two experiments as described in Section III. These experiments aim to address the following questions: (a) Can the Diff-Control policy demonstrate generalization capabilities across diverse tasks? (b) To what extent does the Diff-Control policy outperform the current state-of-the-art methods in terms of overall performance? (c) What are the distinguishing characteristics and benefits of utilizing a stateful policy compared to a non-stateful policy?
We propose the 4 baselines: 1) Image-BC [1], 2) ModAttn [14], 3) BC-Z LSTM, and 4) Diffusion Policy [3]. These baselines all adopt an image-to-action agent framework with varied implementation details. For all experiments, we present the results obtained from the best performing configuration of each baseline method. All the baseline models are reproduced and trained using the collected expert demonstrations for a total of 3,000 epochs. Throughout the training process, checkpoints are saved every 300 epochs. In our analysis, we report the best results achieved from these saved checkpoints for each baseline method.
@@ IV Evaluation
Method Duck Scoop Drum Beats
Scoop Duration Hit $\times$ 3 Duration
Image-BC 16% 21.04$\pm$1.99 0% 19.50$\pm$0.79
ModAttn 12% 20.43$\pm$0.16 0% 25.74$\pm$0.68
BC-Z LSTM 36% 21.21$\pm$1.54 0% 17.53$\pm$0.31
Diffusion Policy 76% 21.72$\pm$2.17 24% 20.60$\pm$1.64
@@
Figure 6: When visual occlusion is applied, Diff-Control (2nd row) manages to scoop the duck from the water successfully. In contrast, the diffusion policy (1st row) fails. Red shade implies perturbation.
IV-A Duck Scooping Evaluation